OpenStudy (anonymous):

Find the value for x, using your knowledge of properties to help you. x*1=5*2

OpenStudy (anonymous):

x = 10

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Welcome ;D

mathslover (mathslover):

no direct answers @SchoolFreak

OpenStudy (anonymous):

lol my bad

mathslover (mathslover):

Well it was said that "using your knowledge" ...

mathslover (mathslover):

@heartbreak try to find : 5*2

mathslover (mathslover):

can you tell me what is 5 * 2 ?

OpenStudy (anonymous):

10

mathslover (mathslover):

good and x *1 = ?

OpenStudy (anonymous):

1x

mathslover (mathslover):

can we remove 1 from there that is x* 1 = x?

mathslover (mathslover):

wrong @SchoolFreak .. x*1=5*2 1x=10 x=10 since x *1 =1x = x

OpenStudy (anonymous):

@mathslover yes

mathslover (mathslover):

hence x = 10 .. (this is the conclusion) @heartbreak try to make it easier by first solving left hand side and then right hand side

OpenStudy (anonymous):